(PA41) The reaction of hydrogen gas with oxygen gas is likely responsible for the Hindenburg disaster...

(PA41)

The reaction of hydrogen gas with oxygen gas is likely responsible for the Hindenburg disaster in 1937, when a airship filled with hydrogen gas ignited and burned, killing 37 people.

If 616 g of oxygen gas reacts completely with hydrogen gas, what volume of water vapor will be produced (in L) at 0.952 atm and 308°C ? Write the balanced equation to help you solve this problem.
Answer #1

Balanced chemical reaction is

2H2 + O2  \rightarrow 2H2O

molar mass of O2 = 31.9988 gm/mole then 616 gm of O2 = 616 / 31.9988 = 19.25 mole

According to reaction 1 mole of O2 produce 2 mole of water then 19.25 mole of O2 produce 19.25 \times 2 / 1 = 38.50 mole of water

to calculate volume of water vapour use ideal gas equation

We know that PV = nRT

V = nRT/P

n = 38.50 mole,

T = 3080C = 308+273.15 = 581.15K,

P= 0.952 atm,

R = 0.08205 L atm mol-1 K-1 ( R = gas constant)

V = ?

Substitute these value in above equation.

V = 38.50 \times 0.08205 \times 581.15 / 0.952 = 1928.37 L

1928.37 L water vapour produced